40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16717     Entry time: Wed Mar 9 10:23:28 2022
Author: Jordan 
Type: Update 
Category: VAC 
Subject: Leak Testing of New Manual Gate Valve - Attempt #1 

Jordan, Chub, Paco

Chub and I went into the lab this morning to leak check the new gate valve after pumping over the weekend. This would be done through the RGA and spraying helium around the newly installed flanges. The RGA is set to monitor the partial pressure of helium versus time and we visually watch for any spikes in pressure to indicate an air leak.

So, I mistakenly thought the gate valve was opened and both sides were being pumped on, this was not the case. The vavle was closed so there was a pressure differential whenI turned the handle it tripped the interlocks and closed V4. I closed VM3 to the RGA volume to prevent the filament from being damaged.

Then the medm screen for the vac controls started flashing rapidly, I closed the window and reopened the controls to find all the panels were white, but we could still see the read only vac screen. So Paco restarted c1vac and the controls were restored. V7 closed, but was then reopened.

We now need to restart pumping with the odd pressure differentials between V4 and VM3.

Plan

- TP2 turned off along with the AUX pump

- VM3 opened to vent RGA volume (RGA turned off, filament cooled)

- Open the manual vent screw on TP1 to bring the RGA+TP1 volume back to atmosphere, now no pressure diff. on V4

- Open V4, restart AUX pump to rough out the volume

- Connect RP1/3 line to the pump scroll, turn on RP1/3 RGA+TP1 volume went to mtorr

- Slowly open the manual gate valve

- Connect AUX pump to TP2 and rough out

- Restart TP2, once at speed open V4, close V6 and turn off RP1/3

Now we are pumping on the RGA/TP1 volume with TP2, leak check attempt #2 will happen tomorrow morning

ELOG V3.1.3-